OpenStudy (anonymous):

Find the first derivative of y=cosx/1+sinx

OpenStudy (anonymous):

-1/(1+sinx)

OpenStudy (anonymous):

\[y = \frac{\cos x}{1 + \sin x}\]use the quotient rule\[u = \cos x\]\[u' = -\sin x\]\[v = 1 + \sin x\]\[v' = \cos x\]\[y' = \frac{u'v - uv'}{v^2}\]

OpenStudy (anonymous):

thanx